Q3b) The demand curve for a consumer is P – d = b/q where d and b are constants. Find the 

price elasticity of demand


1
Expert's answer
2021-04-26T19:30:10-0400

If the demand curve for a consumer is p - d = b/q, where d and b are constants, then:

p = b/q + d and q = b/(p - d).

The price elasticity of demand is:

"Ed = p/q*q' = \frac{-bp} {q×(p - d)^2}."